OpenStudy (anonymous):

What makes the statement true? 8^-2 = ? A. 64 B. 16 C. -1/64 D. 1/64

OpenStudy (anonymous):

D. 8^-2=.015 and if you divide 1/64 it is also .015

OpenStudy (anonymous):

thanks @ohlookidontunderstandanotherthing i understand it now :)

OpenStudy (anonymous):

Yay :D

OpenStudy (anonymous):

8^-2=(8^-1)^2=(1/8)^2=1/64